### v2

@@ 2 Excited states in $d=2$ with geometric modular flows
In the case of an infinite subregion $\mathcal{V}$, we also demand $\xi$ cannot have exotic asymptotic behavior; that is, it must asymptote to the modular flows of either the vacuum state (becoming a vacuum-sector flow) or the thermal state (becoming a thermal-sector flow). [^11] We are glad to thank Jon Sorce for discussions and explanations on this point and its relation to Hilbert space sectors. We refer readers to a forthcoming work by Jon Sorce and Gautam Satishchandran on the topic of boundary conditions for modular flow in free theories.
@@ 2.2 Excited states in the thermal sector
With respect to the Rindler wedge, $|\beta\rangle$ has modular flow Borchers:1998ye
@@ 3.1 Vacuum sector
Moreover, we remark that the null energy densities read
@@ 3.3 An example
Consider the following ansatz for the conformal map:
$u(\tilde{u})=\tilde{u}+e^{-\tilde{u}^{2}}j_{z}(\tilde{u})\qquad v(\tilde{v})=\tilde{v}+e^{-\tilde{v}^{2}}j_{z}(\tilde{v})\,,$ (108)
where the exponential factors are designed to suppress the contribution of $j_{z}$ at large distances, and thus preserve the boost-like property at infinity. By assumption, the Rindler wedges are preserved, so we require $j_{z}(0)=0$. If additionally $j_{z}$ is chosen such that the conformal map is invertible, then the flow described by eq. (107) is an Unruh flow.
For example, let us choose
$j_{z}(\tilde{u}):=z\tanh\tilde{u}\,,$ (109)
with a constant parameter $z$. Over the range $z\in(z_{\min}=-1,z_{\max}\simeq 2.3908)$, $u(\tilde{u})$ is a monotonically increasing function everywhere. At the lower bound $z=-1$, $u(\tilde{u})$ has vanishing first and second derivatives at the origin $\tilde{u}=0$, while at the upper bound $\simeq 2.3908$, similar critical points develop at $\tilde{u}\simeq\pm 1.088$. On the other hand, for $z$ within the range above, the conformal map is invertible everywhere. The corresponding Unruh flow (107), with $z=1.8$, is compared to the vacuum modular flow in figure 3.
@@
Figure 3: Left: vacuum modular flow for $\mathcal{R}$. Right: an Unruh flow described by eqs. (107)-(109) with $z=1.8$. Middle: the difference between the two flows is future-directed in $\mathcal{R}$. By contrast, when $z<0$, the perturbation is past-directed. Future/past directedness of the perturbation is related to the sign of the entanglement entropy difference as seen from eq. (67), see also figure 5.
@@
The energy density (51) is plotted in the left panel of figure 4 for $z=-0.92$ and $2.2$. By design, for $z=0$ we recover the vacuum modular flow which yields vanishing energy density everywhere. The right panel shows the total energy (56) of the corresponding states across the full allowed range of $z$.
@@
Figure 4: Energy density for $z=-0.92$ and $z=2.2$ along the $\tilde{t}=0$ slice (left), and total energy as a function of $z$ (right) for states associated with eqs. (107)-(109).
@@
As anticipated in section 3, the energy density can be negative in some places, but the total energy is always positive. For $z=-0.92\approx z_{\min}$ (blue curve), the total energy becomes large due to the large lumps of positive energy centered at $\tilde{x}=0$, where the conformal map (108) is nearly degenerate. A similar divergence occurs for $z=2.3\approx z_{\max}$ (red curve): the lumps are now located at $\tilde{x}\approx\pm 1.1$, where the conformal map is nearly degenerate. More generally, the sign of $z$ controls whether the energy density has a single set of lumps near $\tilde{x}=0$ ($z<0$ case) or two sets near $\tilde{x}\approx\pm 1.1$ ($z>0$ case). Since from eq. (108) we have $\langle T_{\tilde{u}\tilde{u}}(\tilde{u})\rangle_{\Psi}\big{|}_{\tilde{u}=-\tilde{x}}=\langle T_{\tilde{v}\tilde{v}}(\tilde{v})\rangle_{\Psi}\big{|}_{\tilde{v}=\tilde{x}}=\frac{1}{2}\langle T_{\tilde{t}\tilde{t}}(\tilde{x})\rangle_{\Psi}$, each of the lumps in the energy density shown in figure 4 actually correspond to two overlapping pulses with identical profiles, one traveling left and the other right. The left-travelling pulse eventually escapes from the right Rindler wedge. This does not affect the entanglement entropy, since all Cauchy slices for the Rindler wedge include this pulse. That is, a Rindler observer sees the left-travelling pulse cross the Rindler horizon after infinite proper time.
@@
Figure 5: Left: The orange curve indicates the relative-to-the-vacuum change in the total entanglement entropy of excited states associated with eqs. (107)-(109), while the green curve indicates their relative entropy. Right: Change in the entropy density, i.e., the integrand in eq. (73), for various choices of $z$, including the values $z=-0.92$ and $z=2.2$ considered also in figure 4. An analogous plot for the relative entropy density is provided in figure 6.
@@
The integrand in eq. (73) gives the vacuum-subtracted entropy density for the excited states. This is plotted in the right panel of figure 5 for $z=-0.92$, $-0.7$, $1$, and $2.2$. For $z=0$, the corresponding state is the vacuum and hence the change in the entropy density vanishes everywhere. Compared with the energy density, the simpler profile of the entropy density exhibits a single lump. For $z<0$, the lump is everywhere positive yielding the increased entanglement entropy indicated by the orange curve in the left panel. Similarly, $z>0$ yields a negative lump to produce a decrease in the overall entanglement entropy.
The relative entropy as a function of $z$ is shown in green in the left panel of figure 5. Since the relative entropy is a measure of distinguishability, it vanishes at $z=0$, where |$\ket{\Psi}=\ket{\Omega}$, and grows monotonically as move away from $z=0$, ultimately diverging as $z$ approaches either $z_{\max}$ or $z_{\min}$. This indicates that the states are becoming more and more distinguishable as the absolute value of $z$ increases. While the relative entropy is always positive, the relative entropy density, i.e., the integrand of eq. (80), can be negative, as shown in figure 6. Note again the divergent behavior as $z$ approaches $z_{\min}$ and $z_{\max}$.
Figure 6: Relative entropy density of states associated with eqs. (107-109), for various $z$.
@@ 3.4 Disentangled states
with appropriate sign flips when the arguments are negative. As anticipated, it is singular at the entangling surface, but it can be obtained by taking a limit of regular maps. [^39] For example, consider a map defined by eq. (108) with $j_{z}(\tilde{u}):=\exp\left(-\frac{z}{\tilde{u}^{2}}\right)\tilde{u}^{\frac{1}{\lambda}}$. The perturbation is highly suppressed near the origin for finite $z$, thus preserving the boost-like property, but for $z=0$ ensures a leading $\tilde{u}^{\frac{1}{\lambda}}$ behavior of $u(\tilde{u})$ at the origin. Furthermore, unlike the map discussed below eq. (110), the ansatz (108) ensures the perturbation is suppressed at infinity, thus preserving the vacuum sector property.
@@ 4.2 A state for every flow in the diamond and in the cylinder
As an application of the procedure described above, we can start from the Minkowski vacuum state and create a local excitation such that the resulting state is thermal in the diamond ${\cal D}$. Given a global thermal state, its modular flow with respect to $\mathcal{D}$ Borchers:1998ye ; Calabrese:2004eu yields
@@
Hence, we seek to excite the vacuum such that the resulting state’s modular flow $\xi_{\cal D}$ in $\mathcal{D}$ matches eq. (128). Thanks to theorem 2.2 in Summers:2003tf , the modular flow in a subregion $\mathcal{V}$ uniquely specifies expectation values in $\mathcal{V}$. Therefore, we are demonstrating how an excitation of the Minkowski vacuum can spoof the physics of the thermal state inside $\mathcal{D}$.
@@ 4.3 A unifying picture
While $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ seem to play very different roles in the preceding discussion, we can exchange their roles by flipping the sign convention for $\Omega(x)$, as it is defined in appendix E. Then, $\Omega(x)$ is positive in the preimage of $\mathcal{V}_{2}$ and negative in the preimage of $\mathcal{V}_{1}$. The difference between these two cases can be understood as the difference between sliding the green Rindler wedge in the first figure of 10 upwards until it crosses $\mathcal{I}^{+}$, versus downwards and to the left until it crosses $\mathcal{I}^{-}$. The key point is that nonlocal operators appear whenever a conformal map such as eq. (137) is employed to push subregions beyond infinity, and hence the resulting modular flow is associated with an algebra containing nonlocal operators.
@@ 5.1 Entanglement currents
where $h$ is the determinant of the $H^{d-1}$ metric and $\Omega_{d-2}=\frac{2\pi^{\frac{d-1}{2}}}{\Gamma\left(\frac{d-1}{2}\right)}$ is the area of the unit $(d-2)$-sphere. Moreover, $\alpha_{0}$ is a thermal entropy density; c.p. eqs. (138)-(140) with $\xi=R\partial_{\tau}$ from footnote 49. It can be written as
@@ 6 Discussion
The idea of local entanglement entropy densities has previously been studied independently of geometric modular flows. In particular, Vidal:2014aal advocated for the construction of the “entanglement contour,” i.e., an entropy density $s=s(x)$ whose integral across a given Cauchy surface yields the entanglement entropy for the subregion of interest. The entanglement contour is defined to satisfy several axioms, such as positivity and invariance under the action of a local unitary. This latter requirement distinguishes entanglement contours from our entropy current, as a typical unitary is expected to disrupt the geometric nature of the modular flow, leaving us without a definition of the entropy current, see appendix B. Further, in Vidal:2014aal the contour axioms do not single out a unique $s=s(x)$, while our entropy current is uniquely determined be the modular flow given a state $\Psi$ and a subregion $\mathcal{V}$.
@@
Two sectors at different temperatures are unitarily inequivalent; see footnote 35. Consistency with this statement, in our framework, follows from the requirement that the conformal unitaries act trivially at spatial infinity, such that the vacuum or thermal sector falloffs for the modular flow in eqs. (13) and (40) are preserved. One may wonder how to extend our construction to allow for operators which change the temperature. In this setting, it is important to understand the asymptotic behavior which characterizes the different sectors.
@@
where we specify the curves by where they cross the $x$-axis (i.e., $t=0$) with $v_{0}=x_{0}=-u_{0}$, which occurs at $s=0$. Now consider these curves (179) to be the worldlines describing a family of Unruh observers traversing the diamond. Hence, we may wish to re-express them in terms of the proper time $\tau$ of the observers. This is done using the relation [^61] To derive this, integrate the relation $\frac{\text{d}\tau}{\text{d}s}=\norm{\xi}$ from footnote 29, using eqs. (179) and (153). We have chosen the integration constant such that $\tau=0$ corresponds to $s=0$. Note that in traversing the entire diamond, $-\infty<s<\infty$ while the proper time $\tau$ spans a finite range $-\tau_{\rm max}<\tau<\tau_{\rm max}$ where at its maximum value one has $\cosh\frac{x_{0}\,\tau_{\rm max}}{R^{2}-x_{0}^{2}}=\frac{R}{\sqrt{R^{2}-x_{0}^{2}}}$.
@@
Now given the periodicity in eq. (181) and the symmetry $\Delta s(\tau_{1},\tau_{2})=\Delta s(\tau_{2},\tau_{1})$, the Wightman function then satisfies $W(\tau_{1}+i\beta_{\rm phys},\tau_{2})=W(\tau_{2},\tau_{1})$, which implies a thermal response of the detector at temperature $\beta_{\mathrm{phys}}$, as shown rigorously in Birrell:1982ix ; DeBievre:2006pys . That is, the observer experiences a thermal bath with temperature given by [^63] One may be concerned that the detector will not reach thermal equilibrium because our observers traverse the causal diamond in a finite time — see footnote 61. But the observer can simply initialize the detector with its different energy levels populated in a thermal distribution determined by $\beta_{\rm phys}$. Then, apart from small fluctuations, the observer’s detector remains in equilibrium as she crosses the diamond.
@@
This does not match the physical temperature in eq. (183). Indeed, it assigns a nonzero temperature for the inertial observer with $x_{0}=0$! From this expression, one may also be concerned that the modular temperature will become negative for large values of $\tau$. However, the observers cross the causal diamond in finite proper time — see footnote 61. A more detailed examination shows that $T$ diverges if we follow the observers to either the future or past tips of the causal diamond, but $T$ never becomes negative.
Recall, the modular temperature can also be written as $\beta=2\pi\frac{\text{d}\tau}{\text{d}s}$ (see footnote 29), which is the temperature ascribed to observers in Martinetti:2002sz . [^65] See related discussions in Earman:2011zz . While the authors do not associate a physical meaning to this temperature, they propose it should be related to the fact that observers do not access all the degrees of freedom of the underlying quantum field theory. We see this intuition is correct in that the modular temperature provides a simple prescription (140) to evaluate the entanglement entropy for the region the observers can access.
@@
Nevertheless, our thermodynamics-inspired entropy formula (140) for states with geometric modular flow $\xi$ is on good footing. Recall $\xi$ is a future-directed conformal Killing vector in the subregion of interest, $\mathcal{V}$. Using a singular Weyl transformation, $\xi\big{|}_{\mathcal{V}}$ be transformed into an ordinary Killing field which is future-directed and timelike everywhere; this happened, for example, in section 5.1, when conformally mapping the double cone to a hyperbolic spacetime. In the new conformal frame where $\xi$ is Killing, the modular temperature equals the physical temperature, and this temperature is varying over very large length scales (or not at all, in the double-cone example). Hence, the thermodynamic entropy formula (138) applies. This validates the entanglement entropy formula (140) because by footnote 47, eq. (140) is Weyl-invariant, and hence it is also applicable for the original setting where $\xi$ is only conformally Killing.

### v3

@@ 2 Excited states in $d=2$ with geometric modular flows
In the case of an infinite subregion $\mathcal{V}$, we also demand $\xi$ cannot have exotic asymptotic behavior; that is, it must asymptote to the modular flows of either the vacuum state (becoming a vacuum-sector flow) or the thermal state (becoming a thermal-sector flow). [^11] We are glad to thank Jon Sorce for discussions and explanations on this point and its relation to Hilbert space sectors. We refer readers to a forthcoming work by Jon Sorce and Gautam Satishchandran on the topic of boundary conditions for modular flow in free theories Sorce:2025Upcoming .
@@ 2.2 Excited states in the thermal sector
With respect to the Rindler wedge, $|\beta\rangle$ has modular flow Borchers:1998ye  [^19] The modular flow for a finite interval in a thermal state of a two-dimensional CFT was explored in Mintchev:2022fcp . The Rindler wedge flow (36) is obtained there via the limit $a=0$ and $b\to\infty$.
@@ 3.1 Vacuum sector
Moreover, we remark that the null energy densities read [^28] Note that, due to the $u$ and $v$ dependence of the energy components, we can compute energy densities everywhere — also, say, in the future and past lightcones of the origin — using the following equation.
@@ 3.3 An example
Consider the following ansatz for the conformal map: [^39] Our discussion easily extends to a more general ansatz: $u(\tilde{u})=\tilde{u}+\delta u(\tilde{u}),$ $v(\tilde{v})=\tilde{v}+\delta v(\tilde{v})$ for a wide class of functions $\delta u(\tilde{u})$ and $\delta v(\tilde{v})$ satisfying $\delta u(\tilde{u}=0)=0=\delta v(\tilde{v}=0)$ and $\delta u(\tilde{u}\to\pm\infty)=0=\delta v(\tilde{v}\to\pm\infty)$. Of course, this generalized ansatz introduces right- and left-moving disturbances determined with $\delta u(\tilde{u})$ and $\delta v(\tilde{v})$, respectively.
$u(\tilde{u})=\tilde{u}+\delta u(\tilde{u})=\tilde{u}+z\,\tilde{u}\,e^{-\tilde{u}^{2}}\,,\qquad v(\tilde{v})=\tilde{v}\,,$ (108)
where for simplicity we are only perturbing the coordinate $u\ (=t-x)$. As we will see below this introduces a right-moving excitation over the original vacuum state. The exponential factor in $\delta u$ is designed to suppress the perturbation at large distances, and thus preserve the boost-like property at infinity. Similarly, the factor of $\tilde{u}$ produces boost-like behaviour at the origin and ensures that the Rindler wedges are preserved.
Further, $z$ is a constant controlling the amplitude of the perturbation. We must restrict this amplitude to realize an invertible conformal map with the Unruh flow described by eq. (107). As $z$ approaches $-1$, $u^{\prime}(\tilde{u}=0)$ approaches zero. In fact, with $z=-1$ , the map $u(\tilde{u})$ has vanishing first and second derivatives at $\tilde{u}=0$ and so we must require $z>-1$. Increasing $z$, we find also that $u(\tilde{u})$ has vanishing first derivative at $\tilde{u}\simeq\pm 1.2248$ when $z\simeq 2.2408$. Hence by restricting our attention to $z\in(z_{\min}=-1,z_{\max}\simeq 2.2408)$, we ensure that $u(\tilde{u})$ is an everywhere monotonically increasing function, and thus invertible. The corresponding Unruh flow (107), with $z=2$, is compared to the vacuum modular flow in figure 3. The difference between the two flows shown in the middle panel reveals that our perturbation $\delta u$ creates a right-moving excitation symmetrically arranged about the Rindler horizon at $\tilde{u}=0$. The latter can also be seen by examining the energy density (51) along the $\tilde{x}$-axis, which is plotted in the left panel of figure 4 for $z=-0.9$ and $z=2$.
@@
Figure 3: Left: vacuum modular flow for $\mathcal{R}$. Right: an Unruh flow described by eqs. (107)-(108) with $z=2$. Middle: the difference between the two flows, which is future-directed in $\mathcal{R}$. By contrast, when $z<0$, the perturbation is past-directed. Future/past directedness of the perturbation is related to the sign of the entanglement entropy difference as seen from eq. (67) – see also figure 5. The color of the arrows indicates the intensity of the vector field. Note that along the $\tilde{x}$-axis, the new modular flow becomes nearly null near $\tilde{x}\simeq\pm 1.22$. This behavior arises due to the critical point which emerges at $z=z_{\max}$; we are close to the maximum at $z_{\max}$ with $z=2$ in this plot. Compare figure 4 and caption.
@@
By design, for $z=0$ we recover the vacuum modular flow which yields vanishing energy density everywhere. The right panel shows the total energy (56) of the corresponding states across the full allowed range of $z$.
@@
Figure 4: Energy density for $z=-0.9$ and $z=2$ along the $\tilde{t}=0$ slice (left), and total energy as a function of $z$ (right) for states associated with eqs. (107)–(108). Note that at the critical value $z=z_{\max}$ the energy density is diverging at $\tilde{x}\simeq\pm 1.22$
@@
As anticipated in section 3, the energy density can be negative in some places, but the total energy is always positive. For $z=-0.9\approx z_{\min}$ (blue curve), the total energy becomes large due to the large ridges of positive energy hugging $\tilde{x}=0$, where the conformal map (108) is nearly degenerate. Similarly for $z=2\approx z_{\max}$ (red curve), the total energy is large because there are pairs of large positive-energy ridges now centered about $\tilde{x}\approx\pm 1.22$, where the conformal map is nearly degenerate. More generally, the sign of $z$ controls whether the energy density has a single set of ridges near $\tilde{x}=0$ ($z<0$ case) or two pairs near $\tilde{x}\approx\pm 1.22$ ($z>0$ case). In principle, these profiles for the energy density have contributions coming from both the left-movers and from the right-movers, i.e., $T_{\tilde{t}\tilde{t}}=T_{\tilde{v}\tilde{v}}+T_{\tilde{u}\tilde{u}}$. However, in the present case with eq. (108), we have $\langle T_{\tilde{v}\tilde{v}}(\tilde{v})\rangle_{\Psi}=0$ and $\langle T_{\tilde{u}\tilde{u}}(\tilde{u})\rangle_{\Psi}\neq 0$. Hence the profiles of the energy density shown in figure 4 correspond to pulses traveling to the right, as one can also see directly from the Unruh flow in figure 3.
@@
Figure 5: Left: The orange curve indicates the relative-to-vacuum change in the total entanglement entropy of excited states associated with eqs. (107)-(108), while the green curve indicates their relative entropy. Right: Change in the entropy density, i.e., the integrand in eq. (73), for various choices of $z$, including the values $z=-0.9$ and $z=2$ considered also in figure 4.
@@
The integrand in eq. (73) gives the vacuum-subtracted entropy density for the excited states. This is plotted in the right panel of figure 5 for $z=-0.9$, $-0.7$, $1$, and $2$. For $z=0$, the corresponding state is the vacuum and hence the change in the entropy density vanishes everywhere. Compared with the energy density, the entropy density has a simpler profile which exhibits a single lump. For $z<0$, the lump is everywhere positive, yielding the increased entanglement entropy indicated by the orange curve in the left panel. Further, the peak rises and move to the left to become a divergence at $\tilde{x}=0$ as $z\to z_{\min}$. Similarly, $z>0$ yields a negative lump to produce a decrease in the overall entanglement entropy. As shown in the figure, the profile is much broader in this case and the minimum is only loosely in the vicinity of the critical point $\tilde{x}\simeq 1.22$, even as $z\to z_{\max}$.
The relative entropy as a function of $z$ is shown in green in the left panel of figure 5. Since the relative entropy is a measure of distinguishability, it vanishes at $z=0$, where |$\ket{\Psi}=\ket{\Omega}$, and grows monotonically as move away from $z=0$, ultimately diverging as $z$ approaches either $z_{\max}$ or $z_{\min}$. This indicates that the states are becoming more and more distinguishable as the absolute value of $z$ increases. While the relative entropy is always positive, the relative entropy density, i.e., the integrand of eq. (80), can be negative.
@@ 3.4 Disentangled states
with appropriate sign flips when the arguments are negative. As anticipated, it is singular at the entangling surface, but it can be obtained by taking a limit of regular maps. [^42] For example, consider a map similar to eq. (108), with $u(\tilde{u})=\tilde{u}+e^{-\tilde{u}^{2}-\frac{z}{\tilde{u}^{2}}}\tilde{u}^{\frac{1}{\lambda}}$. The perturbation is highly suppressed near the origin for finite $z$, thus preserving the boost-like property, but for $z=0$ ensures a leading $\tilde{u}^{\frac{1}{\lambda}}$ behavior of $u(\tilde{u})$ at the origin. Furthermore, unlike the map discussed below eq. (109), the ansatz (108) ensures the perturbation is suppressed at infinity, thus preserving the vacuum sector property.
@@ 4.2 A state for every flow in the diamond and in the cylinder
As an application of the procedure described above, we can start from the Minkowski vacuum state and create a local excitation such that the resulting state is thermal in the diamond ${\cal D}$. Given a global thermal state, its modular flow with respect to $\mathcal{D}$ Borchers:1998ye ; Wong:2013gua yields
@@
Hence, we seek to excite the vacuum such that the resulting state’s modular flow $\xi_{\cal D}$ in $\mathcal{D}$ matches eq. (127). Thanks to theorem 2.2 in Summers:2003tf , the modular flow in a subregion $\mathcal{V}$ uniquely specifies expectation values in $\mathcal{V}$. Therefore, we are demonstrating how an excitation of the Minkowski vacuum can spoof the physics of the thermal state inside $\mathcal{D}$ [^46] The ability to spoof various states in a compact subregion via local excitations is closely tied to the mathematical notion of quasiequivalence; see e.g. Verch:1992eg and upcoming work revisiting this topic in Caminiti:JonPaper .
@@ 4.3 A unifying picture
While $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ seem to play very different roles in the preceding discussion, we can exchange their roles by flipping the sign convention for $\Omega(x)$, as it is defined in appendix E. Then, $\Omega(x)$ is positive in the preimage of $\mathcal{V}_{2}$ and negative in the preimage of $\mathcal{V}_{1}$. The difference between these two cases can be understood as the difference between sliding the green Rindler wedge in the first figure of 9 upwards until it crosses $\mathcal{I}^{+}$, versus downwards and to the left until it crosses $\mathcal{I}^{-}$. The key point is that nonlocal operators appear whenever a conformal map such as eq. (136) is employed to push subregions beyond infinity, and hence the resulting modular flow is associated with an algebra containing nonlocal operators.
@@ 5.1 Entanglement currents
where $h$ is the determinant of the $H^{d-1}$ metric and $\Omega_{d-2}=\frac{2\pi^{\frac{d-1}{2}}}{\Gamma\left(\frac{d-1}{2}\right)}$ is the area of the unit $(d-2)$-sphere. Moreover, $\alpha_{0}$ is a thermal entropy density; c.p. eqs. (137)-(139) with $\xi=R\partial_{\tau}$ from footnote 53. It can be written as
@@ 6 Discussion
The idea of local entanglement entropy densities has previously been studied independently of geometric modular flows. In particular, Vidal:2014aal advocated for the construction of the “entanglement contour,” i.e., an entropy density $s=s(x)$ whose integral across a given Cauchy surface yields the entanglement entropy for the subregion of interest. [^63] See related discussions in Botero:2004vpl ; Coser:2017dtb . The entanglement contour is defined to satisfy several axioms, such as positivity and invariance under the action of a local unitary. This latter requirement distinguishes entanglement contours from our entropy current, as a typical unitary is expected to disrupt the geometric nature of the modular flow, leaving us without a definition of the entropy current, see appendix B. Further, in Vidal:2014aal the contour axioms do not single out a unique $s=s(x)$, while our entropy current is uniquely determined be the modular flow given a state $\Psi$ and a subregion $\mathcal{V}$.
@@
Beyond the Modular Wedge
In section 4, we examined Unruh flows in causally complete regions of a conformally flat spacetime and argued that there is always a state in the CFT2 with the corresponding modular flow. One may have noticed that in the figures describing the associated vector field, we have drawn the vectors over the entire spacetime, i.e., beyond the causally complete region of interest, $\mathcal{V}$. That is, we extend the vector field to regions of the spacetime that naively are not accessible by starting in $\mathcal{V}$ and following the modular flow. [^64] We thank the referee for suggesting that we clarify this point.
For example, in figure 1, one might expect the modular flow with respect to the Rindler wedge $\mathcal{R}$ to act only on $\mathcal{R}$ and, by the two-sided property, on its complement $\mathcal{R}^{\prime}$. In this way one would be led to not consider the causal future and past of the origin, i.e., the regions where the associated boost vector (6) is spacelike.
However, on the mathematical side, one can analyze the situation as follows: In a given Hilbert space sector $\mathcal{H}$ of a quantum field theory, to each causally complete region in spacetime one associates a von Neumann algebra of operators contained in $\mathcal{B}(\mathcal{H})$. Once we select a state in $\mathcal{H}$ which is cyclic and separating with respect to, say, the von Neumann algebra of the Rindler wedge, we can construct the associated modular Hamiltonian. The modular Hamiltonian acts on $\mathcal{H}$ and thus can be used to Heisenberg-evolve operators in $\mathcal{B}(\mathcal{H})$. In particular, it can be used to evolve operators that are supported neither in $\mathcal{R}$ nor in its causal complement $\mathcal{R}^{\prime}$. Since, in the specific case of the vacuum, the modular Hamiltonian associated to $\mathcal{R}$ is proportional to the boost generator, when it Heisenberg evolves an operator supported in, say, some subregion in the causal future of the origin, it will evolve it according to the boost vector field in that region.
This discussion can be extended to all of the geometric modular flows constructed in this paper and hence explains the significance of the vector fields shown outside of $\cal{V}$ and its complement $\cal{V}^{\prime}$.
[…]
Let us add here that as noted in footnote 28, the flow determines the energy densities through eq. (57) and these expressions can be applied everywhere throughout the spacetime. In particular, these expressions still apply outside of the regions where the modular flow is timelike.
@@
Two sectors at different temperatures are unitarily inequivalent; see footnote 37. Consistency with this statement, in our framework, follows from the requirement that the conformal unitaries act trivially at spatial infinity, such that the vacuum or thermal sector falloffs for the modular flow in eqs. (13) and (40) are preserved. One may wonder how to extend our construction to allow for operators which change the temperature. In this setting, it is important to understand the asymptotic behavior which characterizes the different sectors.
@@
where we specify the curves by where they cross the $x$-axis (i.e., $t=0$) with $v_{0}=x_{0}=-u_{0}$, which occurs at $s=0$. Now consider these curves (178) to be the worldlines describing a family of Unruh observers traversing the diamond. Hence, we may wish to re-express them in terms of the proper time $\tau$ of the observers. This is done using the relation [^68] To derive this, integrate the relation $\frac{\text{d}\tau}{\text{d}s}=\norm{\xi}$ from footnote 31, using eqs. (178) and (152). We have chosen the integration constant such that $\tau=0$ corresponds to $s=0$. Note that in traversing the entire diamond, $-\infty<s<\infty$ while the proper time $\tau$ spans a finite range $-\tau_{\rm max}<\tau<\tau_{\rm max}$ where at its maximum value one has $\cosh\frac{x_{0}\,\tau_{\rm max}}{R^{2}-x_{0}^{2}}=\frac{R}{\sqrt{R^{2}-x_{0}^{2}}}$.
@@
Now given the periodicity in eq. (180) and the symmetry $\Delta s(\tau_{1},\tau_{2})=\Delta s(\tau_{2},\tau_{1})$, the Wightman function then satisfies $W(\tau_{1}+i\beta_{\rm phys},\tau_{2})=W(\tau_{2},\tau_{1})$, which implies a thermal response of the detector at temperature $\beta_{\mathrm{phys}}$, as shown rigorously in Birrell:1982ix ; DeBievre:2006pys . That is, the observer experiences a thermal bath with temperature given by [^70] One may be concerned that the detector will not reach thermal equilibrium because our observers traverse the causal diamond in a finite time — see footnote 68. But the observer can simply initialize the detector with its different energy levels populated in a thermal distribution determined by $\beta_{\rm phys}$. Then, apart from small fluctuations, the observer’s detector remains in equilibrium as she crosses the diamond.
@@
This does not match the physical temperature in eq. (182). Indeed, it assigns a nonzero temperature for the inertial observer with $x_{0}=0$! From this expression, one may also be concerned that the modular temperature will become negative for large values of $\tau$. However, the observers cross the causal diamond in finite proper time — see footnote 68. A more detailed examination shows that $T$ diverges if we follow the observers to either the future or past tips of the causal diamond, but $T$ never becomes negative.
Recall, the modular temperature can also be written as $\beta=2\pi\frac{\text{d}\tau}{\text{d}s}$ (see footnote 31), which is the temperature ascribed to observers in Martinetti:2002sz . [^72] See related discussions in Earman:2011zz . While the authors do not associate a physical meaning to this temperature, they propose it should be related to the fact that observers do not access all the degrees of freedom of the underlying quantum field theory. We see this intuition is correct in that the modular temperature provides a simple prescription (139) to evaluate the entanglement entropy for the region the observers can access.
@@
Nevertheless, our thermodynamics-inspired entropy formula (139) for states with geometric modular flow $\xi$ is on good footing. Recall $\xi$ is a future-directed conformal Killing vector in the subregion of interest, $\mathcal{V}$. Using a singular Weyl transformation, $\xi\big{|}_{\mathcal{V}}$ be transformed into an ordinary Killing field which is future-directed and timelike everywhere; this happened, for example, in section 5.1, when conformally mapping the double cone to a hyperbolic spacetime. In the new conformal frame where $\xi$ is Killing, the modular temperature equals the physical temperature, and this temperature is varying over very large length scales (or not at all, in the double-cone example). Hence, the thermodynamic entropy formula (137) applies. This validates the entanglement entropy formula (139) because by footnote 51, eq. (139) is Weyl-invariant, and hence it is also applicable for the original setting where $\xi$ is only conformally Killing.